MASS-SELECTIVE MATRIX-ISOLATION SPECTROSCOPY: FT-IR SPECTRA OF CYANOACETYLENE AND TETRACYANOETHYLENE IONS AND THEIR IONIC FRAGMENTS

\underline{A. M. SMITH-GICKLHORN}, M. FRANKOWSKI, M. LORENZ, V. E. BONDYBEY, Institut für Physikalische und Theoretische Chemie, Lichtenbergstraße 4, D-85747 Garching, Germany; and R. KOŁOS, Institute of Physical Chemistry of the Polish Academy of Sciences, Kasprzaka 44, 01-224 Warsaw, Poland.

Recent spectroscopic results from a new setup for the deposition of mass-selected ions in neon matrices are presented. Gaseous precursor substances are ionized in a homebuilt electron impact source, the ions extracted and deflected by 90^{o} into a quadrupole mass filter with unity mass resolution up to 500 amu. The mass-selected cations are deposited simultaneously with an excess of electrons, to maintain overall sample neutrality and prevent buildup of space charges, and neon gas onto a reflective substrate held at 7 K. Fourier-transform infrared spectra from such samples using cyanoacetylene and tetracyanoethylene reveal new infrared transitions for several cations. The strong infrared signal observed for the asymmetric stretch of CNC^{+} and the appearance of the corresponding band for ^{13}CNC^{+} in its natural isotopic abundance, demonstrates the sensitivity and usefulness of the apparatus. A combination band of CNC^{+} is also observed, revealing its low-frequency bend. Using this setup, the infrared spectra of cyanoacetylene ions^{a} (H(D)C_{3}N^{+} and C_{3}N_{2}^{+}) and tetracyanoethylene ions and ionic fragments^{b} (C_{6}N_{4}^{+}, C_{6}N_{4}^{−}, C_{6}N_{3}^{+}, C_{5}N_{3}^{+}, and C_{3}N_{2}^{+}) are observed for the first time. The assignment of the numberous observed vibrations of these ions are well-supported by our density-functional calculations. Using this apparatus and the sensitive LIF techique, the assignment of a rich emission spectrum with origin near 440 nm is confirmed to be due to mass 88 or C_{5}N_{2} neutral ^{c} as subsequently verified by a cavity ringdown measurement of Linnartz, Maier et al.

aA.M. Smith-Gicklhorn, M. Lorenz, R. Kołos and V.E. Bondybey J. Chem. Phys. \underline{115}, 7534-7542 (2001).

bA.M. Smith-Gicklhorn, M. Frankowski, and V.E. Bondybey Phys. Chem. Chem. Phys. In Press.

cA.M. Smith-Gicklhorn, M. Lorenz, M. Frankowski, R. Kołos and V.E. Bondybey Chem. Phys. Lett. \underline{351}, 85-91 (2002).